# Gaussian is All You Need: A Unified Framework for Solving Inverse Problems via Diffusion Posterior Sampling

### v2

@@ 1 Introduction
Diffusion-based models [17, 34, 38, 39] have recently gained attention due to their powerful generative ability by learning complex data distributions. A number of recent methods have used diffusion-based models as priors for solving inverse problems [21, 48]. Diffusion Posterior Sampling (DPS) [9] and $\Pi$GDM [37] are notable examples that incorporate widely adopted posterior sampling schemes in the reverse diffusion process. Nevertheless, both methods are computationally inefficient as they require computing gradients through the diffusion model to compute the conditional score. Furthermore, DPS uses inexact likelihood covariance matrix estimates (even for simple Gaussian priors).
@@
In this paper, we show that the Gaussian prior assumption is all that is needed to achieve the best of both worlds, offering both covariance correction and computational efficiency. Building on this assumption, we propose a novel and simple sampling framework that is not only computationally efficient but also achieves performance that is competitive with, and in some cases surpasses, existing methods. Our work begins by computing the distribution of measurements conditioned on intermediate outputs of the diffusion network for simple Gaussian priors. For a linear inverse problem with a Gaussian prior, the conditional distribution has an explicit Gaussian form. This distribution has the same form as the conditional distribution proposed in $\Pi$GDM, even though both methods start from different assumptions (see Sec. 3.1). Furthermore, for the isotropic covariance case, our conditional distribution is equivalent to the one used in DPS, up to a scalar constant (see Sec. 3.1). While the conditional distribution is generally intractable for non-Gaussian priors, the form obtained under the Gaussian assumption serves as an approximation that results in remarkable reconstruction performance. Figure 1 presents some examples that illustrate how our method successfully recovers fine details that other methods fail to restore. Another key advantage of this assumption is that it allows us to bypass gradient computation through the network when computing the score of the conditional distribution. Our experiments show that bypassing gradients through the network can provide at least 2×2$\times$2 × reduction in the computational cost.
The Gaussian assumption for approximating the conditional score is already used, both implicitly and explicitly, in different diffusion-based posterior sampling methods. For instance, in $\Pi$GDM the conditional Gaussian assumption can be shown to be equivalent to the Gaussian prior assumption. As a consequence of this assumption, the method was able to introduce a covariance term in the likelihood. A recent method, DDS [10], introduces a novel sampling strategy that entirely bypasses gradient computations through the network. Instead, it approximates the gradient by projecting orthogonally onto the clean data manifold. This approach significantly simplifies the sampling process by assuming that the prior distribution is Gaussian with infinite variance.
@@ 2.2 Posterior sampling for inverse problems
where $\bm{y}$ represents measurements, $\mathcal{A}(\cdot)$ represents forward operator or a measurement matrix, and $\sigma_{n}$ represents noise. The maximum a posteriori (MAP) ${\bm{x}}_{\text{MAP}}$ estimate can be found by maximizing the sum of the likelihood $p(\bm{y}|\bm{x}_{0})$ and the data prior distribution $p(\bm{x}_{0})$. From equation (5), it is evident that $p(\bm{y}|\bm{x}_{0})$ is a Gaussian distribution. However, identifying a suitable prior $p(\bm{x}_{0})$ remains an active area of research. Pre-trained Diffusion models are suitable candidates as they learn the log prior density $\nabla_{\bm{x}}\log p(\bm{x})$. To utilize them for solving inverse problems, we build upon the unconditional denoising prediction in (3) to develop a measurement-conditioned sampling procedure. This involves incorporating a conditional score in the reverse sampling process as
@@
The term $\nabla_{\bm{x}_{t}}\log p(\bm{x}_{t})$ in (6) can be approximated using a pre-trained diffusion model $\bm{s}_{\theta}(\bm{x}_{t},t)$. However, as noted in [37, 9], $p(\bm{y}|\bm{x}_{t})$ is intractable. This issue is discussed in DPS [9] and $\Pi$GDM [37], where it is highlighted that there is no analytical form available for $p(\bm{y}|\bm{x}_{t})$. To understand why this is the case, let us consider
@@
where $\hat{\bm{x}}_{0}$ is given by (3) and $p(\bm{y}|\hat{\bm{x}}_{0})\simeq\mathcal{N}(\mathcal{A}(\hat{\bm{x}}_{0}),\sigma^{2}_{n}\bm{I})$. Alternatively, $\Pi$GDM [37] assumes $p(\bm{x}_{0}|\bm{x}_{t})\sim\mathcal{N}(\hat{\bm{x}}_{0},r_{t}^{2}\bm{I})$, where $r_{t}$ is a time- and data-dependent hyper-parameter, and approximates the conditional score as
@@
where $\bm{\Sigma}_{t}=r_{t}^{2}\bm{A}\bm{A}^{\mathsf{H}}+\sigma_{n}^{2}\bm{I}$ Both approximations in DPS and $\Pi$GDM have demonstrated remarkable results on various real-world datasets. However, the approximation in DPS may become inaccurate as $t$ approaches the total number of diffusion steps $T$. Additionally, both methods compute backpropagation through the diffusion network to evaluate $\partial\hat{\bm{x}}_{0}/\partial\bm{x}_{t}$, which can significantly increase the time and memory complexity of the sampling process.
@@
Several recent works have proposed strategies to eliminate the expensive gradient computation through the diffusion network during inverse problem solving. One such approach is DMPS [24], which introduces a closed-form approximation of the conditional likelihood by assuming an uninformative prior on $p(\bm{x}_{0})$. This assumption allows the method to avoid backpropagation entirely. Notably, our Gaussian prior also represents DMPS as a special case when $\sigma_{0}^{2}\to\infty$. DDNM [41] introduces a zero-shot framework for solving linear inverse problems by leveraging the structure of the forward operator. It decomposes the solution into two orthogonal components: the row space and the null space of the operator $\bm{A}$. The row-space component is obtained directly via the pseudoinverse projection $\bm{A}^{\dagger}\bm{y}$, ensuring strict data consistency. The null-space component is refined by iteratively denoising with a pretrained diffusion model and projecting the result onto the null space using $(\mathbf{I}-\bm{A}^{\dagger}\bm{A})\hat{\bm{x}}_{0}$. However, this method is unstable in the presence of measurement noise. To address this, the authors proposed DDNM+, which includes a range-space correction term. In our experiments, we observed poor performance of DDNM+ on deblurring tasks in the presence of noise, despite the correction term. ReSample [35] is another method that leverages pretrained latent diffusion models (LDMs). It avoids gradient computation by enforcing hard data consistency in the latent space through constrained optimization, a process referred to as a ReSample timestep. During the early stages of the reverse sampling process, the method performs unconditional DDIM updates without applying any ReSample steps. Resampling occurs every $k$ iterations, where $k$ is typically set to 10. Each ReSample step involves two stages: first, minimizing the reconstruction error $\|\bm{y}-\bm{A}(\mathcal{D}(\bm{z}))\|_{2}^{2}$, where $\bm{z}$ is the latent variable initialized from $\bm{\hat{z}}_{0}$ (the latent-space equivalent of the predicted mean in Eq. (3)) and $\mathcal{D}$ is the decoder that maps latent vectors back to the image space. Once the minimizer is obtained, the second stage performs stochastic resampling to generate the new latent variable $\bm{z}_{t}$. Despite only applying sampling at certain steps, ReSample has significantly higher runtime. Additionally, it requires both an encoder and a decoder to map data to and from the latent space, which adds complexity and limits its applicability in settings without pretrained auto-encoder models. Other representative method is DDRM [21], which performs spectral-domain posterior sampling but relies on SVD of the forward operator and assumes linearity. DiffPIR [48] follows a plug-and-play strategy, applying pretrained diffusion models as proximal denoisers without explicitly modeling the posterior. Recent extensions [28] propose covariance estimation modules to enhance conditioning accuracy in such frameworks.
@@
A complementary line of work, referred to as variational inference-based methods, approximates the posterior distribution with a simpler, tractable distribution instead of seeking for a closed-form expression for the conditional score  [11]. RED-Diff [23] proposes to minimize the KL divergence between the true conditional posterior $p(\bm{x}_{0}\mid\bm{y})$ and a variational distribution $q(\bm{x}_{0}|\bm{y})=\mathcal{N}(\bm{\mu},\sigma^{2}\bm{I})$. The mean $\bm{\mu}$ is iteratively updated by minimizing the KL objective, while the variance of the variational distribution is assumed to be small $\sigma^{2}\approx 0$. While RED-Diff and CoDPS share the goal of modeling the posterior that utilize simple Gaussian forms, the two approaches differ significantly in formulation and implementation. RED-Diff treats posterior sampling as a stochastic optimization problem and relies on iterative updates guided by the denoiser, whereas CoDPS derives a closed-form conditional score correction under a Gaussian prior.
Other approaches, such as [8, 39, 7, 22, 25, 20], utilize pre-trained diffusion models. In this work, we demonstrate the connections between DPS, $\Pi$GDM, and DDS-type conditioning methods, and propose a unified diffusion posterior sampling algorithm. Our approach leverages implicit assumptions present in previous works and enhances computational efficiency by eliminating gradient computation through the network. While the focus of our work is mainly on unsupervised diffusion based methods, several deep-learning based inverse problem in a fully-supervised manner [33, 31, 43, 44, 12, 45].
@@ 3.1 Conditional score approximation
We revisit the conditional score approximation from DPS [9], stated as $p(\bm{y}|\bm{x}_{t})\approx\mathcal{N}(\bm{A}\hat{\bm{x}}_{0},\sigma_{n}^{2}\bm{I})$ for linear inverse problem under additive Gaussian measurement noise, where $\hat{\bm{x}}_{0}$ is the quantity given in (3). Intuitively, this approximation appears to be valid for small values of $t$ as the diffusion model learns the true data distribution $p(\bm{x}_{0})$ and the clean image estimates $\hat{\bm{x}}_{0}$ will accurately represent samples from this distribution. However, at the early stages of the reverse diffusion process $t\approx T$, the samples $\bm{x}_{t}$ will be close to an isotropic Gaussian distribution. The clean image estimates $\hat{\bm{x}}_{0}$ from $\bm{x}_{t}$ at these stages will not necessarily come from the true prior distribution. This can be easily verified by looking at the estimates $\hat{\bm{x}}_{0}$ for different values of $t$, as shown in Figure 2. When $1000600<t<$ 1000, the estimates do not seem to resemble clean natural images (i.e samples from $p(\bm{x}_{0})$). Additionally, within the range $t\in[200,600]$, although the samples appear plausible, they are perceptually different from the ground truth image. More importantly, the uncertainty in the measurement $\bm{y}$ when conditioned on these estimates grows as $t$ increases and needs to be accounted for in the conditional variance. To see this more clearly, let us consider a simple prior Gaussian distribution given by $\bm{x}_{0}\sim\mathcal{N}(0,\sigma_{0}^{2}\bm{I})$. Using the properties of marginal and conditional Gaussian distributions, we obtain
@@
In contrast, $\Pi$GDM [37] assumes $p(\bm{x}_{0}|\bm{x}_{t})\sim\mathcal{N}(\hat{\bm{x}}_{0},r_{t}^{2}\bm{I})$, which leads to the approximation in (9). This approximation is equivalent to DPS up to a scalar constant for cases where $\sigma_{n}^{2}\gg r_{t}$, or for the cases where the forward operator has orthogonal rows ( i.e $\bm{A}\bm{A^{\mathsf{H}}}=\bm{I}$). Furthermore, by applying Theorem 1, we can draw a direct connection between the underlying assumptions made $\Pi$GDM and DDS. In particular, we show that the conditional assumption in $\Pi$GDM is equivalent to the Gaussian assumption in DDS.
@@ 3.2 Covariance corrected posterior sampling
0:  $N$, $\bm{y}$, $\zeta_{t}\}_{t=1}^{N},{\{\tilde{\sigma}_{t}\}_{t=1}^{N}}$, $\sigma_{0}^{2}$
1:  $\bm{x}_{N}\sim\mathcal{N}(\bm{0},\bm{I})$
2:  for $i=N-1$ to 00 do
3:     $t,\bar{\alpha}_{t},\bar{\alpha}_{t-1}\leftarrow\text{extract\_alpha\_t}(i)$triangleright Get alpha values
@@
7:     $\bar{\bm{x}}_{t}\leftarrow\sqrt{\bar{\alpha}_{t-1}}\hat{\bm{x}}_{0}+c_{1}\bm{z}+c_{2}\hat{\bm{s}}$triangleright DDIM coefficients
8:      $\bm{x}_{t-1}\leftarrow\bar{\bm{x}}_{t}+\zeta_{t}\bm{\kappa}_{t}$triangleright Data consistency update (12)
@@
Our proposed score function in (12) is directly obtained from (13) (see the derivation in Appendix A.3). While $\bm{\kappa}_{t}$ in (12) shares certain similarities with (9) proposed by $\Pi$GDM, a key distinction is that our method avoids the computationally intensive Jacobian term ${\partial\hat{\bm{x}}_{0}}/{\partial\bm{x}_{t}}$. We avoid this expensive computation as a direct consequence of our assumption on $p(\bm{x}_{0})$, as shown in the following derivation:
@@
=$\stackrel{{\scriptstyle(i)}}{{=}}\frac{1}{{\sqrt{\bar{\alpha}}}_{t}}+\frac{(1-{\bar{\alpha}}_{t})}{{\sqrt{\bar{\alpha}}_{t}}}\dfrac{\partial^{2}}{\partial\bm{x}_{t}^{2}}\log p(\bm{x}_{t})$
=$\stackrel{{\scriptstyle(ii)}}{{=}}\frac{1}{{\sqrt{\bar{\alpha}}}_{t}}-\dfrac{1-{\bar{\alpha}}_{t}}{{\sqrt{\bar{\alpha}}_{t}}(1-\bar{\alpha}_{t}+\sigma_{0}^{2}\bar{\alpha}_{t})}.$ (14)
@@
Deblur (Gaussian) Deblur (motion) SR (×$\times 4$)
@@
DDS [10] 100 27.14 0.777 49.73 0.300 20.48 0.542 87.53 0.443 25.63 0.709 66.36 0.373
ReSample [35] 100 25.14 0.615 57.92 0.378 23.75 0.513 0.697 0.474 19.33 0.292 143.4 0.627
$\Pi$GDM [37] 100 20.02 0.447 104.81 0.560 27.41 0.782 35.00 0.251 23.79 0.598 69.82 0.392
DMPS [24] 1000 25.55 0.724 28.76 0.259 - - - - 26.33 0.754 26.55 0.247
DPS [9] 1000 26.12 0.748 26.26 0.237 23.96 0.678 29.82 0.286 25.08 0.710 29.52 0.270
DiffPIR [48] 100 24.58 0.674 29.87 0.298 26.92 0.757 25.38 0.254 22.96 0.666 46.83 0.357
DDRM [21] 20 25.90 0.741 57.88 0.303 - - - - 26.47 0.761 56.41 0.299
MCG [8] 1000 6.72 0.051 101.2 0.340 6.72 0.055 310.5 0.702 20.05 0.559 87.64 0.520
PnP-ADMM [6] - 23.58 0.684 94.25 0.418 23.43 0.669 87.23 0.450 22.14 0.592 135.24 0.530
ADMM-TV [15] - 24.63 0.721 101.83 0.460 21.95 0.655 158.43 0.519 21.14 0.618 258.06 0.580

### v3

@@ 1 Introduction
[…]
@@
[…]
@@ 2.2 Posterior sampling for inverse problems
where $\bm{y}$ represents measurements, $\mathcal{A}(\cdot)$ represents a forward operator or a measurement matrix, and $\sigma_{n}$ represents noise. The maximum a posteriori (MAP) ${\bm{x}}_{\text{MAP}}$ estimate can be found by maximizing the sum of the log likelihood $\log p(\bm{y}|\bm{x}_{0})$ and the log of the data prior distribution $\log p(\bm{x}_{0})$. From equation (5), it is evident that $p(\bm{y}|\bm{x}_{0})$ is a Gaussian distribution. However, identifying a suitable prior $p(\bm{x}_{0})$ remains an active area of research. Pre-trained Diffusion models are suitable candidates as they learn the log prior density $\nabla_{\bm{x}}\log p(\bm{x})$. To utilize them for solving inverse problems, we build upon the unconditional denoising prediction in (3) to develop a measurement-conditioned sampling procedure. This involves incorporating a conditional score in the reverse sampling process as
@@
$p(\bm{y}|\bm{x}_{t})=\int p(\bm{y}|\bm{x}_{0},\bm{x}_{t})p(\bm{x}_{0}|\bm{x}_{t})\,d\bm{x}_{0}.$ (7)
Since $\bm{x}_{t}$ is obtained from $\bm{x}_{0}$ by corrupting it with noise, once we know $\bm{x}_{0}$, knowing $\bm{x}_{t}$ does not provide any additional information about $\bm{y}$. For this reason, we can write the conditional as $p(\bm{y}|\bm{x}_{t})=\int_{\bm{x}_{0}}p(\bm{y}|\bm{x}_{0})p(\bm{x}_{0}|\bm{x}_{t})\,d\bm{x}_{0}.$ The main challenge then lies in modeling $p(\bm{x}_{0}|\bm{x}_{t})$. Using the reverse diffusion sampling process, we can approximate and sample from this distribution. However, marginalizing over $\bm{x}_{0}$ is infeasible, as generating even a single sample requires performing a complete reverse ancestral sampling. To circumvent this, DPS[7] proposed to approximate the conditional score $\bm{\kappa}_{t}$ as
@@
[…]
@@
Several recent works have proposed strategies to eliminate the expensive gradient computation through the diffusion network during inverse problem solving. One such approach is DMPS [14], which introduces a closed-form approximation of the conditional likelihood by assuming an uninformative prior on $p(\bm{x}_{0})$. This assumption allows the method to avoid backpropagation entirely. Notably, our Gaussian prior also represents DMPS as a special case when $\sigma_{0}^{2}\to\infty$. DDNM [15] introduces a zero-shot framework for solving linear inverse problems by leveraging the structure of the forward operator. It decomposes the solution into two orthogonal components: the row space and the null space of the operator $\bm{A}$. The row-space component is obtained directly via the pseudoinverse projection $\bm{A}^{\dagger}\bm{y}$, ensuring strict data consistency. The null-space component is refined by iteratively denoising with a pretrained diffusion model and projecting the result onto the null space using $(\mathbf{I}-\bm{A}^{\dagger}\bm{A})\hat{\bm{x}}_{0}$. However, this method is unstable in the presence of measurement noise. To address this, the authors proposed DDNM+, which includes a range-space correction term. In our experiments, we observed poor performance of DDNM+ on deblurring tasks in the presence of noise, despite the correction term. ReSample [16] is another method that leverages pretrained latent diffusion models (LDMs). It avoids gradient computation by enforcing hard data consistency in the latent space through constrained optimization, a process referred to as a ReSample timestep. ReSample applies unconditional DDIM updates in the early stages, and resampling is performed every $k$ steps (typically $k=10$). Despite only applying sampling at certain steps, ReSample has significantly higher runtime. Additionally, it requires both an encoder and a decoder to map data to and from the latent space, which adds complexity and limits its applicability in settings without pretrained auto-encoder models.
DiffPIR [6] follows a plug-and-play strategy and uses pretrained diffusion models as generative denoiser priors. While DiffPIR does not explicitly model the conditional posterior, the sampling steps in DiffPIR closely resemble the steps in DPS and CoDPS (albeit in a different order, as a result of different assumptions). Our experimental results show that CoDPS provides better and improved results compared to DPS and DiffPIR. Ather representative method is DDRM [5], which performs spectral-domain posterior sampling but relies on SVD of the forward operator and assumes linearity. DiffPIR [6] follows a plug-and-play strategy, applying pretrained diffusion models as proximal denoisers without explicitly modeling the posterior. Recent extensions [17] propose covariance estimation modules to enhance conditioning accuracy in such frameworks.
@@
A complementary line of work, referred to as variational inference-based methods, approximates the posterior distribution with a simpler, tractable distribution instead of seeking for a closed-form expression for the conditional score  [18]. RED-Diff [19] proposes to minimize the KL divergence between the true conditional posterior $p(\bm{x}_{0}\mid\bm{y})$ and a variational distribution $q(\bm{x}_{0}|\bm{y})=\mathcal{N}(\bm{\mu},\sigma^{2}\bm{I})$. The mean $\bm{\mu}$ is iteratively updated by minimizing the KL objective, while the variance of the variational distribution is assumed to be small $\sigma^{2}\approx 0$. While RED-Diff and CoDPS share the goal of modeling the posterior that utilize simple Gaussian forms, the two approaches differ significantly in formulation and implementation. RED-Diff formulates posterior sampling as a stochastic optimization problem that requires iterative optimization with reverse diffusion sampling in each step. In contrast, CoDPS derives a closed-form correction to the score under a Gaussian prior without any need to perform an iterative optimization problem.
[…]
@@ 3.1 Conditional score approximation
We revisit the conditional score approximation from DPS [7], stated as $p(\bm{y}|\bm{x}_{t})\approx\mathcal{N}(\bm{A}\hat{\bm{x}}_{0},\sigma_{n}^{2}\bm{I})$ for a linear inverse problem under additive Gaussian measurement noise, where $\hat{\bm{x}}_{0}$ is the quantity given in (3). Intuitively, this approximation appears to be valid for small values of $t$ as the diffusion model learns the true data distribution $p(\bm{x}_{0})$ and the clean image estimates $\hat{\bm{x}}_{0}$ will accurately represent samples from this distribution. However, at the early stages of the reverse diffusion process $t\approx T$, the samples $\bm{x}_{t}$ will be close to an isotropic Gaussian distribution. The clean image estimates $\hat{\bm{x}}_{0}$ from $\bm{x}_{t}$ at these stages will not necessarily come from the true prior distribution. This can be easily verified by looking at the estimates $\hat{\bm{x}}_{0}$ for different values of $t$, as shown in Figure 2. When $600<t<1000$, the estimates do not seem to resemble clean natural images (i.e samples from $p(\bm{x}_{0})$). Additionally, within the range $t\in[200,600]$, although the samples appear plausible, they are perceptually different from the ground truth image. More importantly, the uncertainty in the measurement $\bm{y}$ when conditioned on these estimates grows as $t$ increases and needs to be accounted for in the conditional variance. To see this more clearly, let us consider a simple prior Gaussian distribution given by $\bm{x}_{0}\sim\mathcal{N}(0,\sigma_{0}^{2}\bm{I})$. Using the properties of marginal and conditional Gaussian distributions, we obtain
@@
[…]
@@ 3.2 Covariance corrected posterior sampling
0:NN, $\bm{y}$, $\zeta_{t}\}_{t=1}^{N},{\{\tilde{\sigma}_{t}\}_{t=1}^{N}}$, $\sigma_{0}^{2}$
[…]
2:for i=N1i=N-1 to 0 do
3:  $t,\bar{\alpha}_{t},\bar{\alpha}_{t-1}\leftarrow\text{extract\_alpha\_t}(i)$ $\triangleright$ Get alpha values
@@
7:  $\bar{\bm{x}}_{t}\leftarrow\sqrt{\bar{\alpha}_{t-1}}\hat{\bm{x}}_{0}+c_{1}\bm{z}+c_{2}\hat{\bm{s}} \triangleright$ DDIM coefficients
8:   $\bm{x}_{t-1}\leftarrow\bar{\bm{x}}_{t}+\zeta_{t}\bm{\kappa}_{t}$ $\triangleright$ Data consistency update (12)
@@
[…]
@@
$\stackrel{{\scriptstyle(i)}}{{=}}\frac{1}{{\sqrt{\bar{\alpha}}}_{t}}+\frac{(1-{\bar{\alpha}}_{t})}{{\sqrt{\bar{\alpha}}_{t}}}\dfrac{\partial^{2}}{\partial\bm{x}_{t}^{2}}\log p(\bm{x}_{t})$
$\stackrel{{\scriptstyle(ii)}}{{=}}\frac{1}{{\sqrt{\bar{\alpha}}}_{t}}-\dfrac{1-{\bar{\alpha}}_{t}}{{\sqrt{\bar{\alpha}}_{t}}(1-\bar{\alpha}_{t}+\sigma_{0}^{2}\bar{\alpha}_{t})}.$ (14)
@@
Deblur (Gaussian) Deblur (motion) SR ($\times 4$)
@@
DDS [9]
100 27.14 0.777 49.73 0.300 20.48 0.542 87.53 0.443 25.63 0.709 66.36 0.373
ReSample [16]
100 25.14 0.615 57.92 0.378 23.75 0.513 0.697 0.474 19.33 0.292 143.4 0.627
[…]
100 20.02 0.447 104.81 0.560 27.41 0.782 35.00 0.251 23.79 0.598 69.82 0.392
DMPS [14]
1000 25.55 0.724 28.76 0.259 - - - - 26.33 0.754 26.55 0.247
DPS [7]
1000 26.12 0.748 26.26 0.237 23.96 0.678 29.82 0.286 25.08 0.710 29.52 0.270
DiffPIR [6]
100 24.58 0.674 29.87 0.298 26.92 0.757 25.38 0.254 22.96 0.666 46.83 0.357
DDRM [5]
20 25.90 0.741 57.88 0.303 - - - - 26.47 0.761 56.41 0.299
MCG [20]
1000 6.72 0.051 101.2 0.340 6.72 0.055 310.5 0.702 20.05 0.559 87.64 0.520
PnP-ADMM [33]
- 23.58 0.684 94.25 0.418 23.43 0.669 87.23 0.450 22.14 0.592 135.24 0.530
ADMM-TV [34]
- 24.63 0.721 101.83 0.460 21.95 0.655 158.43 0.519 21.14 0.618 258.06 0.580